# Identifying Majorana edge and end modes in Josephson junction of $p$-wave superconductor with magnetic barrier

### v1

@@ I Introduction
Majorana fermions are the particles of their own antiparticle that manifest as zero-energy quasiparticles in topological superconductors  Kitaev (2001); Alicea (2012); Leijnse and Flensberg (2012). These modes are distinguished by their unique non-Abelian exchange statistics, allowing them for the manipulation of quantum information through braiding operations Ivanov (2001); Nayak et al. (2008). Moreover, their intrinsic robustness against local perturbations makes them highly attractive for fault-tolerant topological quantum computation  Kitaev (2001); Ivanov (2001); Nayak et al. (2008); Kitaev (2009); Alicea (2012); Leijnse and Flensberg (2012); Beenakker (2013), as they are less susceptible to decoherence and environmental noise. In current literature, a one-dimensional (1D) $p$-wave superconductors (SCs) provide a fundamental platform for the realization of end localized Majorana zero modes (MZMs), as first demonstrated by Kitaev in his seminal work Kitaev (2001).
@@
In order to engineer such $p$-wave pairing starting from a conventional $s$-wave SC, several theoretical proposals exists in literature. The first elegant proposal is based on 1D Rashba nanowire, where a semiconducting nanowire with strong Rashba spin-orbit coupling (SOC) in presence of an applied Zeeman field is proximity coupled to a conventional $s$-wave SC Oreg et al. (2010); Lutchyn et al. (2010); Mondal et al. (2023). Very recently, altermagnets Šmejkal et al. (2022a, b) have been proposed to realize MZMs replacing the external Zeeman field Ghorashi et al. (2024); Mondal et al. (2025). These setups provide a compelling indirect experimental signatures of MZMs through the zero bias peak in differential conductance measurements Oreg et al. (2010); Lutchyn et al. (2010); Mondal et al. (2025); Das et al. (2012); Mourik et al. (2012). Beyond Rashba nanowire model, another insightful proposal is utilized in a system consisting of an array of magnetic adatoms deposited on top of a conventional $s$-wave SC in both 1D and 2D domain  Pientka et al. (2013); Nadj-Perge et al. (2013); Klinovaja et al. (2013); Braunecker and Simon (2013); Vazifeh and Franz (2013); Sau and Demler (2013); Pientka et al. (2014); Pöyhönen et al. (2014); Reis et al. (2014); Hu et al. (2015); Hui et al. (2015); Hoffman et al. (2016); Christensen et al. (2016); Sharma and Tewari (2016); Andolina and Simon (2017); Kaladzhyan et al. (2017); Theiler et al. (2019); Sticlet and Morari (2019); Mashkoori and Black-Schaffer (2019); Ménard et al. (2019); Mashkoori et al. (2020); Teixeira et al. (2020); Rex et al. (2020); Perrin et al. (2021); Kobiałka et al. (2020); Chatterjee et al. (2023a, b); Mondal et al. (2023); Yang et al. (2016); Pöyhönen et al. (2016). In this scenario, the interplay between magnetic impurity spins and the $s$-wave SC gives rise to in-gap Yu-Shiba-Rusinov (YSR) bound states Pientka et al. (2013); Yazdani (2015); Shiba (1968). The hybridization of YSR states form Shiba band, which effectively mimicks as a $p$-wave SC in the minigap. These engineered heterostructures successfully create the necessary conditions to support topological superconductivity hosting localized MZMs. Later such advancements have significantly propelled the experimental study of Majorana quasiparticles in magnet-SC heterostructures Yazdani et al. (1997, 1999); Yazdani (2015); Schneider et al. (2021); Beck et al. (2021); Wang et al. (2021); Schneider et al. (2022); Küster et al. (2022); Lo Conte et al. (2022); Yazdani et al. (2023); Soldini et al. (2023).
@@
Josephson junctions (JJs) are among the most extensively studied systems in condensed matter physics not only because they exhibit intriguing physical phenomena but also due to their promising applications Josephson (1962); Golubov et al. (2004); Makhlin et al. (2001); Bergeret et al. (2005); Birge and Satchell (2024). Furthermore, recent advancements in JJs engineered in hybrid superconductor systems have opened new avenues for realizing and manipulating Majorana modes Rokhinson et al. (2012); Suleymanli et al. (2020); Nakhmedov et al. (2020); Abboud et al. (2022); Kwon et al. (2004); Fu and Kane (2008); Potter and Fu (2013); Hegde et al. (2020). In this regard, JJs composed of $p$-wave SCs on both sides of a nonmagnetic potential barrier under external magnetic field have been studied Nakhmedov et al. (2020); Abboud et al. (2022). Additionally, JJs involving unconventional SCs hybridised with topological insulators have also been investigated Ye et al. (2019); Williams et al. (2012). Very recently, an alternative approach to external magnetic field is utilized by placing magnetic spin textures at the junction between two conventional $s$-wave SCs to induce an effective magnetic field and SOC Sardinero et al. (2024). This effectively generates MZMs localized at the two ends of the junction. However, efffects of such magnetic barrier placed at the junction between 2D $p$-wave SCs are not explored yet. In general, it is very interesting to separate the edge localized Majorana modes and MZMs localized at the end of the magnetic barrier via any suitable response.
@@
Figure 1: Schematic diagram of our JJ setup is depicted which comprises of two 2D $p$-wave SCs with opposite phase differences of ±$\pm\phi/2$ in the left and right regions. The junction is characterized by tunneling through a spatially modulated magnetic barrier. MEMs arise in the intrinsic topological $p$-wave SCs, while localized MZMs are generated at two ends of the interfacial magnetic barrier.
@@
Motivated by this fact, in this article, we introduce a theoretical model of a JJ composed of 2D $p$-wave SCs featuring a magnetically textured barrier, as illustrated in Fig. 1. We investigate the impact of both $p_{x}+p_{y}$ and $p_{x}+ip_{y}$ pairing symmetries incorporated in the SC and ask the following intriguing questions: (a) how does the strength and spatial periodicity of the magnetic barrier influence the system’s topological properties? (b) how can the Josephson current (JC) possibly identify different Majorana modes present in the system, even if they are hybridized? We intend to answer these questions in this article. We uncover three distinct topological phases revealing the localization of 1D Majorana edge modes at the edges of the SC and MZMs at the ends of the magnetic barrier. Additionally, we analyze the hybridization of these modes via local density of states (LDOS). Furthermore, we also compute the JC associated with these three phases and identify the signatures of each phases. By exploring the interplay between magnetic textures and superconducting pairing symmetries, our findings provide insights about the Majorana modes that appear in this hybrid setup.
@@
The remainder of the paper is organized as follows. In Sec. II, we introduce our Hamiltonian of the 2D tight-binding model describing two $p$-wave SCs and a magnetic textured barrier at the junction. In Sec. III, we discuss the effect of barrier on the band spectrum and the LDOS. Sec. IV is devoted to the analysis of the impact of the phase difference on LDOS. We discuss the identification of Majorana modes in different topological phases via the JC signal in Sec. V. Finally, we summarize our findings and conclude in Sec. VI.
@@ II Model
$\mathcal{H}_{L/R}=-t\!\!\!\!\sum_{<\mathbf{r},\mathbf{r}^{\prime}>,\sigma}\!\!\!\!c^{\dagger}_{\mathbf{r},\sigma}\,c_{\mathbf{r}^{\prime},\sigma}\,\!\!\!$ $+$ $\!\!\!\!\sum_{<\mathbf{r},\mathbf{r}^{\prime}>,\sigma}\!\!\!\Delta_{\mathbf{r},\mathbf{r}^{\prime}}e^{i\phi_{L/R}}\,c^{\dagger}_{\mathbf{r},\sigma}c^{\dagger}_{\mathbf{r}^{\prime},\sigma}+\rm{h.c}$ (2)
-- $\mu\sum_{\mathbf{r},\sigma}c^{\dagger}_{\mathbf{r},\sigma}\,c_{\mathbf{r},\sigma}\ ,$
where, $c_{\mathbf{r},\sigma}$ and $c_{\mathbf{r},\sigma}^{\dagger}$ denote the annihilation and creation operators for electrons with spin $\sigma$ at site $\mathbf{r}=(x,y)$. Model parameters $t$ and $\mu$ dentoe the hopping amplitude and chemical potential of the two $p$-wave SCs respectively. The superconducting phase difference between the right and left SCs is $\phi=\phi_{R}-\phi_{L}$. Importantly, the superconducting pairing amplitude, $\Delta_{\mathbf{r},\mathbf{r}^{\prime}}$ is different for $(p_{x}+p_{y})$- and $(p_{x}+ip_{y})$ pairing symmetry and given as, Zhang et al. (2019); Wang et al. (2017); Röntynen and Ojanen (2015); Zhou et al. (2014)
$\Delta_{\mathbf{r},\mathbf{r}^{\prime}}=\!\!(\Delta_{x,x+a},\Delta_{y,y+a})=\begin{cases}(\Delta_{0},\Delta_{0})\mathrm{\,\,for\,\,}(p_{x}+p_{y})\mathrm{\,\,pairing}\\ (\Delta_{0},i\Delta_{0})\mathrm{\,\,for\,\,}(p_{x}+ip_{y})\mathrm{\,\,pairing}\\ \end{cases}$
Here, $\Delta_{0}$ denotes the bulk superconducting pairing amplitude of the considered $p$-wave SCs.
@@
where, $\tilde{t}_{y}^{\sigma\sigma^{\prime}}$ represents the tunneling matrix element of the following $2\times 2$ matrix in spin space, incorporating the magnetic texture barrier and is given by,
@@ III Effect of Magnetic Barrier
$N(\mathbf{r},\omega)=-\frac{1}{\pi}\mathrm{Im[Tr}\,\mathcal{G(\mathbf{r},\mathbf{r},\omega)}]\ ,$ (7)
where, $\mathcal{G}(\mathbf{r},\mathbf{r^{\prime}},\omega)$ is the retarded Green’s function and obtain using Eq. (1) as,
$\mathcal{G}(\mathbf{r},\mathbf{r^{\prime}},\omega)=[(\omega+i\delta)\mathbf{I}-\mathcal{H}_{\rm{JJ}}(\mathbf{r},\mathbf{r^{\prime}})]^{-1}\ .$ (8)
Here, we mainly focus on the zero energy LDOS i.e. $N(\mathbf{r},\omega=0)$ in the $x-y$ plane to characterize and understand the localization profile of MEMs as well as MZMs. We set $t=1,~{}\Delta_{0}=0.5t,~{}t_{0}=1$, and $L_{x}=L_{y}=50$ lattice sites thoughout the paper. Other model parameters values are mentioned explicitly. The results of this analysis are presented and discussed below.
Figure 2: We illustrate the energy eigenvalue spectrum in the absence ($t_{m}=0$) and presence ($t_{m}\neq 0$) of a magnetic barrier for a $p_{x}+p_{y}$ type SC in panels (a) and (b) respectively. Panels (c) and (d) depict the LDOS for $\mu=3t$ and $\mu=4.2t$ under the influence of the magnetic barrier ($t_{m}\neq 0$). For panels (b), (c), and (d), the barrier strength is chosen as $t_{m}=2t$. The calculations are performed using OBC and on a finite size $50\times 50$ square lattice. The other model parameters are chosen as $\Delta_{0}=0.5t$, $\xi_{m}=5$, $t_{0}=t$, and $t=1$.
@@ III.1 For $p_{x}+p_{y}$ type pairing
A ($p_{x}\!+\!p_{y}$)-wave SC harbors intrinsic gapless topological superconductivity hosting MFEMs at zero energy Zhang et al. (2019); Wang et al. (2017). In Fig. 2(a), we plot the energy eigenvalue spectra of Hamiltonian [Eq.(1)] as a function of chemical potential, $\mu$. We clearly observe the existence of zero energy MFEMs for $\mu<4t$ showcasing the topology of 2D ($p_{x}+p_{y}$)-wave SC. For $\mu>4t$, the bulk of the system becomes gapped and MFEMs also disappear from the system. In absence of the magnetic barrier ($t_{m}=0$) only the spin-conserving hopping parameter, $t_{0}$, in Eq. (4) contributes. In this case, the two SCs behave as a single ($p_{x}+p_{y}$)-wave SC, and consequently, zero-energy modes localize at the edges, corresponding to MFEMs which is schematically illustrated in Fig. 1. Now, when the strength of the magnetic barrier is turned on, two other topological phases emerge in the system. In Fig. 2(b), we illustrate the eigenvalues spectra for $t_{m}=2t,~{}\xi_{m}=5$ as a funciton of $\mu$. First. we observe that for $\mu_{c1}<\mu<4t$, the system hosts both MZMs and MFEMs. Note, for $t_{m}=2t$, we find $\mu_{c1}\simeq 1.64t$ which, in general, will depend on the values of $t_{m}$ and $\xi_{m}$. Presence of such mixed topological phase, is further confirmed by computing the LDOS, $N(\mathbf{r},E=0)$. We plot the $N(\mathbf{r},E=0)$ in the $x-y$ plane for $\mu=3t$ in Fig. 2(c) and observe that the MZMs are localized at the ends of the magnetic barrier whereas MFEMs are localized along the edges of the 2D domain. Note that, the junction MFEMs originating from each of the isolated $p_{x}+p_{y}$-type SC are now gapped out due to the coupling Hamiltonian, $\mathcal{H}_{LR}$ [see Eq. (3)]. Furthermore, from Fig. 2(b) we observe the presence of only MZMs beyond $\mu=4t$ confined solely to the end of the magnetic barrier, while the MFEMs of the $p$-wave SC are completely gapped out. This is further assured from LDOS calculation as shown in Fig. 2(d) where we plot the $N(\mathbf{r},E=0)$ for $\mu=4.2t$. These observations suggests the appearance of two new topological phases and more importantly extension of topological regime beyond $\mu=4t$ compared to bare $p$-wave SC. We enlist these three topological phases below,

### v2

# Identifying Majorana edge and end modes in a Josephson junction of a $p$-wave superconductor with a magnetic barrier

@@ I Introduction
Majorana fermions are the particles of their own antiparticle that manifest as zero-energy quasiparticles in topological superconductors  Kitaev (2001); Alicea (2012); Leijnse and Flensberg (2012). These modes are distinguished by their unique non-Abelian exchange statistics, allowing them for the manipulation of quantum information through braiding operations Ivanov (2001); Nayak et al. (2008). Moreover, their intrinsic robustness against local perturbations makes them highly attractive for fault-tolerant topological quantum computation  Kitaev (2001); Ivanov (2001); Nayak et al. (2008); Kitaev (2009); Alicea (2012); Leijnse and Flensberg (2012); Beenakker (2013); Tanaka et al. (2024, 2012), as they are less susceptible to decoherence and environmental noise. In current literature, a one-dimensional (1D) $p$-wave superconductors (SCs) provide a fundamental platform for the realization of end localized Majorana zero modes (MZMs), as first demonstrated by Kitaev in his seminal work Kitaev (2001).
@@
In order to engineer such $p$-wave pairing starting from a conventional $s$-wave SC, several theoretical proposals exists in literature. The first elegant proposal is based on 1D Rashba nanowire, where a semiconducting nanowire with strong Rashba spin-orbit coupling (SOC) in presence of an applied Zeeman field is proximity coupled to a conventional $s$-wave SC Oreg et al. (2010); Lutchyn et al. (2010); Mondal et al. (2023a). Very recently, altermagnets Šmejkal et al. (2022a, b) have been proposed to realize MZMs replacing the external Zeeman field Ghorashi et al. (2024); Mondal et al. (2025). These setups provide a compelling indirect experimental signatures of MZMs through the zero bias peak in differential conductance measurements Oreg et al. (2010); Lutchyn et al. (2010); Mondal et al. (2025); Das et al. (2012); Mourik et al. (2012). Beyond Rashba nanowire model, another insightful proposal is utilized in a system consisting of an array of magnetic adatoms deposited on top of a conventional $s$-wave SC in both 1D and 2D domain  Pientka et al. (2013); Nadj-Perge et al. (2013); Klinovaja et al. (2013); Braunecker and Simon (2013); Vazifeh and Franz (2013); Sau and Demler (2013); Pientka et al. (2014); Pöyhönen et al. (2014); Reis et al. (2014); Hu et al. (2015); Hui et al. (2015); Hoffman et al. (2016); Christensen et al. (2016); Sharma and Tewari (2016); Andolina and Simon (2017); Kaladzhyan et al. (2017); Theiler et al. (2019); Sticlet and Morari (2019); Mashkoori and Black-Schaffer (2019); Ménard et al. (2019); Mashkoori et al. (2020); Teixeira et al. (2020); Rex et al. (2020); Perrin et al. (2021); Kobiałka et al. (2020); Chatterjee et al. (2023a, b); Mondal et al. (2023a); Yang et al. (2016); Pöyhönen et al. (2016). In this scenario, the interplay between magnetic impurity spins and the $s$-wave SC gives rise to in-gap Yu-Shiba-Rusinov (YSR) bound states Pientka et al. (2013); Yazdani (2015); Shiba (1968). The hybridization of YSR states form Shiba band, which effectively mimicks as a $p$-wave SC in the minigap. These engineered heterostructures successfully create the necessary conditions to support topological superconductivity hosting localized MZMs. Later such advancements have significantly propelled the experimental study of Majorana quasiparticles in magnet-SC heterostructures Yazdani et al. (1997, 1999); Yazdani (2015); Schneider et al. (2021); Beck et al. (2021); Wang et al. (2021); Schneider et al. (2022); Küster et al. (2022); Lo Conte et al. (2022); Yazdani et al. (2023); Soldini et al. (2023).
@@
Josephson junctions (JJs) are among the most extensively studied systems in condensed matter physics not only because they exhibit intriguing physical phenomena but also due to their promising applications Josephson (1962); Golubov et al. (2004); Makhlin et al. (2001); Bergeret et al. (2005); Birge and Satchell (2024). Furthermore, recent advancements in JJs engineered in hybrid superconductor systems have opened new avenues for realizing and manipulating Majorana modes Rokhinson et al. (2012); Suleymanli et al. (2020); Nakhmedov et al. (2020); Abboud et al. (2022); Kwon et al. (2004); Fu and Kane (2008); Potter and Fu (2013); Hegde et al. (2020); Pal et al. (2025). In this regard, JJs composed of $p$-wave SCs in presence of a diffusive region and on both sides of a nonmagnetic potential barrier under external magnetic field have been studied Nakhmedov et al. (2020); Abboud et al. (2022); Asano et al. (2006). Additionally, JJs involving unconventional SCs hybridised with topological insulators have also been investigated Ye et al. (2019); Williams et al. (2012). Very recently, an alternative approach to external magnetic field is utilized by placing magnetic spin textures at the junction between two conventional $s$-wave SCs to induce an effective magnetic field and SOC Sardinero et al. (2024). This effectively generates MZMs localized at the two ends of the junction. However, efffects of such magnetic barrier placed at the junction between 2D $p$-wave SCs are not explored yet. In general, it is very interesting to separate the edge localized Majorana modes and MZMs localized at the end of the magnetic barrier via any suitable response.
Recent advancements have underscored the potential of 2D platforms in realizing and simulating quantum gates through braiding operations. Notably, simulations of quantum gates using MZMs localized in magnetic vortices of 2D topological superconductors have been successfully demonstrated Buss et al. (2025). Furthermore, hybrid architectures combining MEMs and MZMs within vortex cores offer a promising route toward efficient and fault-tolerant quantum gate implementation Bedow and Morr (2025). These developments emphasize the practical relevance of 2D systems for topological quantum computation and provide strong motivation for our current investigation. In particular, a key advantage of 2D models lies in their suitability for braiding unlike 1D systems that rely on complex configurations such as T-junctions to exchange MZMs Alicea (2012), whereas 2D geometries inherently support adiabatic exchange and direct braiding of MZMs in real space.
@@
Figure 1: Schematic diagram of our $J$ setup is depicted which comprises of two 2D $p$-wave SCs with opposite phase differences of $\pm\phi/2$ in the left and right regions. The junction is characterized by tunneling through a spatially modulated magnetic barrier. MEMs arise in the intrinsic topological $p$-wave SCs, while localized MZMs are generated at two ends of the interfacial magnetic barrier.
@@
Motivated by this fact, in this article, we introduce a theoretical model of a JJ composed of 2D $p$-wave SCs featuring a magnetically textured barrier, as illustrated in Fig. 1. We investigate the impact of both $p_{x}+p_{y}$ and $p_{x}+ip_{y}$ pairing symmetries incorporated in the SC and ask the following intriguing questions: (a) how does the strength and spatial periodicity of the magnetic barrier influence the system’s topological properties? (b) how can the Josephson current (JC) possibly identify different Majorana modes present in the system, even if they are hybridized? We intend to answer these questions in this article. We uncover three distinct topological phases revealing the localization of 1D Majorana edge modes at the edges of the SC and MZMs at the ends of the magnetic barrier. Additionally, we analyze the hybridization of these modes via local density of states (LDOS). Furthermore, we also compute the JC associated with these three phases and identify the signatures of each phases. In particular, the Josephson current exhibits a discontinuous jump due to the edge modes and pronounced hump in the $p_{x}+ip_{y}$ pairing case, directly linked to the hybridized Majorana modes.
@@
The remainder of the paper is organized as follows. In Sec. II, we introduce our Hamiltonian of the 2D tight-binding model describing two $p$-wave SCs and a magnetic textured barrier at the junction. In Sec. III, we discuss the effect of barrier on the band spectrum and the LDOS. Sec. IV is devoted to the analysis of the impact of the phase difference on LDOS. We discuss the identification of Majorana modes in different topological phases via the JC signal in Sec. V. We analyze the stability of MZMs and the JC in Sec. VI by introducing a perturbation at the end of the magnetic barrier coupling to the SCs. Finally, we summarize our findings and conclude in Sec. VII. The derivation of the effective 1D Hamiltonian and the topological characterization are provided in Appendix A and Appendix B, respectively. In Appendix C, we also discuss the relevance of our results to the corresponding JC behavior of 1D $p$-wave Kitaev chains with a magnetic weak link.
@@ II Model
[…]
where, $c_{\mathbf{r},\sigma}$ and $c_{\mathbf{r},\sigma}^{\text{\textdagger}}$ denote the annihilation and creation operators for electrons with spin $\sigma$ at site $\mathbf{r}=(x,y)$. Model parameters $t$ and $\mu$ denote the hopping amplitude and chemical potential of the two $p$-wave SCs respectively. The superconducting phase difference between the right and left SCs is $\phi=\phi_{R}-\phi_{L}$. Importantly, the superconducting pairing amplitude, $\prime\Delta_{\mathbf{r},\mathbf{r}^{\prime}}$ is different for $(p_{x}+p_{y})$- and $(p_{x}+ip_{y})$ pairing symmetry and given as, Zhang et al. (2019); Wang et al. (2017); Röntynen and Ojanen (2015); Zhou et al. (2014)
[…]
Here, $\Delta_{0}$ denotes the bulk superconducting pairing amplitude of the considered $p$-wave SCs. Note that, $(p_{x}+p_{y})$ and $(p_{x}+ip_{y})$ pairing exhibit distinct topological band properties as we discuss later.
@@
where, $\prime\tilde{t}_{y}^{\sigma\sigma^{\prime}}$ represents the tunneling matrix element of the following $2\times 2$ matrix in spin space, incorporating the magnetic texture barrier and is given by,
@@ III Effect of Magnetic Barrier
[…]
Here, we mainly focus on the zero energy LDOS i.e. $N(\mathbf{r},\omega=0)$ in the xyx-y plane to characterize and understand the localization profile of MEMs as well as MZMs. We set $t=1,~\Delta_{0}=0.5t,~t_{0}=1$, and $L_{x}=L_{y}=50$ lattice sites thoughout the paper. Other model parameters values are mentioned explicitly. The results of this analysis are presented and discussed below.
Figure 2: We illustrate the energy eigenvalue spectrum in the absence ($t_{m}=0$) and presence ($t_{m}\ne 0$) of a magnetic barrier for a $p_{x}+p_{y}$ type SC in panels (a) and (b) respectively. The insets of panel (a) and (b) exhibits the zoomed version of eigenvalue spectrum near zero-energy for better clarity. Panels (c) and (d) depict the LDOS for $\mu=3t$ and $\mu=4.2t$ under the influence of the magnetic barrier ($t_{m}\ne 0$). For panels (b), (c), and (d), the barrier strength is chosen as $t_{m}=2t$. The calculations are performed using OBC and on a finite size $50\times 50$ square lattice. The other model parameters are chosen as $\Delta_{0}=0.5t$, $\xi_{m}=5$, $t_{0}=t$, and $t=1$.
@@ III.1 For $p_{x}+p_{y}$ type pairing
A ($p_{x}\!+\!p_{y}$)-wave SC harbors intrinsic gapless topological superconductivity hosting MFEMs at zero energy Zhang et al. (2019); Wang et al. (2017). In Fig. 2(a), we plot the energy eigenvalue spectra of Hamiltonian [Eq.(1)] as a function of chemical potential, $\mu$. We clearly observe the existence of zero energy MFEMs for $\mu<4t$ showcasing the topology of 2D ($p_{x}+p_{y}$)-wave SC. For $\mu>4t$, the bulk of the system becomes gapped and MFEMs also disappear from the system. In absence of the magnetic barrier ($t_{m}=0$) only the spin-conserving hopping parameter, $t_{0}$, in Eq. (4) contributes. In this case, the two SCs behave as a single ($p_{x}+p_{y}$)-wave SC, and consequently, zero-energy modes localize at the edges, corresponding to MFEMs which is schematically illustrated in Fig. 1. Now, when the strength of the magnetic barrier is turned on, two other topological phases emerge in the system. In Fig. 2(b), we illustrate the eigenvalues spectra for $t_{m}=2t,~\xi_{m}=5$ as a funciton of $\mu$. First. we observe that for $\mu_{c1}<\mu<4t$, the system hosts both MZMs and MFEMs. Note, for $t_{m}=2t$, we find $\mu_{c1}\simeq 1.64t$ which, in general, will depend on the values of $t_{m}$ and $\xi_{m}$. Presence of such mixed topological phase, is further confirmed by computing the LDOS, $N(\mathbf{r},E=0)$. We plot the $N(\mathbf{r},E=0)$ in the xyx-y plane for $\mu=3t$ in Fig. 2(c) and observe that the MZMs are localized at the ends of the magnetic barrier whereas MFEMs are localized along the edges of the 2D domain. Note that, the junction MFEMs originating from each of the isolated $p_{x}+p_{y}$-type SC are now gapped out due to the coupling Hamiltonian, $\mathcal{H}_{LR}$ [see Eq. (3)]. Furthermore, from Fig. 2(b) we observe the presence of only MZMs beyond $\mu=4t$ confined solely to the end of the magnetic barrier, while the MFEMs of the $p$-wave SC are completely gapped out. This is further assured from LDOS calculation as shown in Fig. 2(d) where we plot the $N(\mathbf{r},E=0)$ for $\mu=4.2t$. These observations suggests the appearance of two new topological phases and more importantly extension of topological regime beyond $\mu=4t$ compared to bare $p$-wave SC. We enlist these three topological phases below,
@@
For the purpose of topological characterization of these phases hosting different modes, we derive an effective 1D Hamiltonian by setting $L_{x}=2$ at the junction and considering the limit $L_{y}\rightarrow\infty$ (see Appendix A for details), thereby reducing the full 2D system to two superconducting chains coupled along the $y$-direction via the magnetic barrier. For the resulting effective Hamiltonian $H_{\text{eff}}$ (see Appendix B.2), we compute the topological invariant, the winding number Chiu et al. (2016); Mondal et al. (2023b) in the $t_{m}$-$\xi_{m}$ parameter space. This winding number characterizes the presence of MZMs at the end points of the junction barrier. The invariant $W$ takes values 0, 1, or 2, corresponding respectively to: (i) a trivial phase with no MZMs, (ii) the presence of two MZMs (one at each end of the junction), and (iii) four MZMs (two at each end) as discussed in Appendix A.